OpenStudy (anonymous):

Write the quadratic function f(x) = x^2 - 5x + 3 in vertex form. A) f(x) = (x - 2.5)^2 + 3 B) f(x) = (x + 2.5)^2 + 3 C) f(x) = (x - 2.5)^2 - 3.25 D) f(x) = (x + 2.5)^2 - 3.25

OpenStudy (anonymous):

plz let me know how to solve this type of problem too plz

OpenStudy (anonymous):

first coordinate of the vertex is found via \[-\frac{b}{2a}\] in your case it is \[-\frac{-5}{2\times 1}=\frac{5}{2}=2.5\] so you know it will look like \[y=(x-2.5)^2+k\]

OpenStudy (anonymous):

so find \(k\) notice that if you replace \(x\) by \(2.5\) you get \[y=(2.5-25)^2+k=0+k=k\] so replace \(x\) by \(2.5\) in the original expression \[y=x^2-5x+3\] and see what you get. that will be \(k\)

OpenStudy (anonymous):

k hold on

OpenStudy (anonymous):

-3.25

OpenStudy (anonymous):

=k

OpenStudy (anonymous):

right

OpenStudy (anonymous):

Its C thank you!

OpenStudy (anonymous):

yw